# Low-Dimension-to-High-Dimension Generalization And Its Implications for Length Generalization

### v1

@@ A.3 Proof for Theorem 5
$\dot{A}=-\frac{1}{D_{N_{0}}}\sum_{x\in{\mathcal{X}}_{N}}\left(\left\langle xe_{n(x)}^{\intercal},A\right\rangle-c^{*}(x)\right)\sum_{k=1}^{r}\left\langle xe_{n(x)}^{\intercal},U_{k}\right\rangle U_{k}.$
@@
=$=\begin{cases}-\frac{1-x_{j}}{2}\prod_{k=j+1}^{N}\frac{1+x_{k}}{2},&i=j,\\ x_{i}\frac{1-x_{j}}{2}\prod_{k=j+1}^{N}\frac{1+x_{k}}{2},&i<j.\end{cases}$
@@
  • If $b_{k}(x)=0$ for all $x\in{\mathcal{X}}_{0}$, we have $\hat{p}_{k}=0$ and $p_{k}^{2}\geq 0=\hat{p}_{k}^{2}$;
  • If $b_{k}(x)\neq 0$ for some $x\in{\mathcal{X}}_{0}$, we have $\hat{p}_{k}=p_{k}$ and thus $\hat{p}_{k}^{2}=p_{k}^{2}$. To show this, without loss of generality, we suppose that for $1\leq k\leq m$, there exists some $x\in{\mathcal{X}}_{N_{0}}$ such that $b_{k}(x)\neq 0$, and for $m+1\leq k\leq r$, $b_{k}(x)=0$ for all $x\in{\mathcal{X}}_{N_{0}}$. It suffices to show that the following equation has a unique solution:
@@
A.4 Proof for Corollary 3
By Theorem 5, the PLAA with GRPE achieves LHDH generalization from ${\mathcal{X}}_{N_{0}}$ to ${\mathcal{X}}_{N}$ if and only if the target concept $c(x)$ is the min-degree interpolator w.r.t. the independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$. Therefore, it is equivalent to prove the target concept $c(x)$ is the min-degree interpolator w.r.t. the independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$ if and only if
@@ A.4 Proof for Corollary 3
We first show the sufficiency. Assume that $c(x)$ is not the min-degree interpolator w.r.t. the independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$ on ${\mathcal{X}}_{N_{0}}$. In other words, there exists an interpolator $\tilde{c}(x)=\sum_{k=1}^{r}\tilde{c}_{k}\sum_{1\leq i\leq j\leq N}(U_{k})_{ij}b_{ij}^{\text{PLAA}}(x)$ on ${\mathcal{X}}_{N_{0}}$ such that
@@
$\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(c)\leq\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(\tilde{c}),$
@@
which contradicts that the target concept $c(x)$ is the min-degree interpolator w.r.t. the independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$ on ${\mathcal{X}}_{N_{0}}$.
Appendix B Experiments
B.1 Position Embeddings for Length Generalization
For the URF $n$-addition training data, we first sample the lengths of two addends uniformly from $\{1,\dots,n\}\times\{1,\dots,n\}$. For two addends of lengths $(n_{1},n_{2})$, we then samples from $[10^{n_{1}-1},10^{n_{1}}-1]\times[10^{n_{2}-1},10^{n_{2}}-1]$ (If $n_{i}=0$, then the corresponding sample interval is $[0,9]$). This is to guarantee the addends are length-uniform. For the addition $x_{n_{1}-1}\dots x_{0}+y_{n_{2}-1}\dots y_{0}=z_{n_{3}}z_{n_{3}-1}\dots z_{0}$, the training instance is
@@ B.1 Position Embeddings for Length Generalization
Here, we add spaces between the characters to ensure each is tokenized separately. We use “$\mathtt{b}$” and “$\mathtt{e}$” instead of “$\mathtt{[BOS]}$” and “$\mathtt{[EOS]}$” for a simpler implementation with the GPT-2 tokenizer. In our experiments, we train with $10000$ URF $4$-addition samples.
We choose GPT-2 with key-only position embeddings as our model. For the RPE and RPE-Square settings, we augment the GPT-2 model with RPE and RPE-Square, respectively. The implementation is adapted from HuggingFace [23].
We train the models by AdamW, with the initial learning rate $0.0005$, the weight decay $1.0$, and the cosine scheduler. The warmup ratio is $0.05$. The gradient accumulation steps is $2$. The per-device training batch size is $128$. We train the models for $20000$ steps and $200000$ steps. The experiments are run on a server with Ubuntu. All the models are trained on two NVIDIA GeForce RTX 3090 GPUs.

### v3

@@ C.5 Proof for Theorem 4
$\dot{A}=-\frac{1}{D_{N_{0}}}\sum_{x\in{\mathcal{X}}_{N_{0}}}\left(\left\langle xe_{n(x)}^{\intercal},A\right\rangle-c^{*}(x)\right)\sum_{k=1}^{r}\left\langle xe_{n(x)}^{\intercal},U_{k}\right\rangle U_{k}.$
@@
[…]
@@
  • If $b_{k}(x)=0$ for all $x\in{\mathcal{X}}_{N_{0}}$, we have $\hat{p}_{k}=0$ and $p_{k}^{2}\geq 0=\hat{p}_{k}^{2}$;
  • If $b_{k}(x)\neq 0$ for some $x\in{\mathcal{X}}_{N_{0}}$, we have $\hat{p}_{k}=p_{k}$ and thus $\hat{p}_{k}^{2}=p_{k}^{2}$. To show this, without loss of generality, we suppose that for $1\leq k\leq m$, there exists some $x\in{\mathcal{X}}_{N_{0}}$ such that $b_{k}(x)\neq 0$, and for $m+1\leq k\leq r$, $b_{k}(x)=0$ for all $x\in{\mathcal{X}}_{N_{0}}$. It suffices to show that the following equation has a unique solution:
@@
C.6 Proof for Corollary 2
By Theorem 4, the PLAA with GRPE achieves LHDH generalization from ${\mathcal{X}}_{N_{0}}$ to ${\mathcal{X}}_{N}$ if and only if the target concept $c(x)$ is the min-degree interpolator w.r.t. the linearly independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$. Therefore, it is equivalent to prove the target concept $c(x)$ is the min-degree interpolator w.r.t. the linearly independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$ if and only if
@@ C.6 Proof for Corollary 2
We first show the sufficiency. Assume that $c(x)$ is not the min-degree interpolator w.r.t. the linearly independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$ on ${\mathcal{X}}_{N_{0}}$. In other words, there exists an interpolator $\tilde{c}(x)=\sum_{k=1}^{r}\tilde{c}_{k}\sum_{1\leq i\leq j\leq N}(U_{k})_{ij}b_{ij}^{\text{PLAA}}(x)$ on ${\mathcal{X}}_{N_{0}}$ such that
@@
[…]
@@
which contradicts that the target concept $c(x)$ is the min-degree interpolator w.r.t. the linearly independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$ on ${\mathcal{X}}_{N_{0}}$.
Appendix D Experiments
D.1 Detailed Explanation of RPE-Square
The design of RPE-Square follows the principle: when devising position embeddings for length generalization, one needs to consider both the LDHD generalization in the latent space and the nuisance of the data format in the input sequence space. This principle is derived from the LDHD generalization perspective for length generalization, illustrating the insights of LDHD generalization for practical models.
To further elaborate, we show how RPE-Square is constructed in the guidance of the proposed principle to achieve length generalization of the URF addition (with CoT).
  • LDHD generalization in the latent space. For the addition, the LDHD generalization in the latent space can be handled by RPE.
[…]
    $s=\begin{cases}x_{n_{1}(s)-1},&n_{1}(s)>0,n_{2}(s)<0,n_{3}(s)<0,\\ y_{n_{2}(s)-1},&n_{1}(s)>0,n_{2}(s)>0,n_{3}(s)<0,\\ z_{n_{3}(s)-1},&n_{1}(s)>0,n_{2}(s)>0,n_{3}(s)>0.\end{cases}$
    Note that $z_{k}$ can be determined by $x_{k-1}$, $x_{k}$, $y_{k-1}$, $y_{k}$, and $z_{k}$ (we ignore the boundary cases in this discussion for simplicity). To predict the next token of some $s$ where $n_{1}(s)>0,n_{2}(s)>0,n_{3}(s)>0$, we need to identify the elements $s_{1},s_{2},s_{3},s_{4},s_{5}$ such that
    $\left\{\begin{matrix}n_{1}(s_{1})=n_{3}(s)-1,&n_{2}(s_{1})<0,&n_{3}(s_{1})<0,\\ n_{1}(s_{2})=n_{3}(s),&n_{2}(s_{2})<0,&n_{3}(s_{2})<0,\\ n_{2}(s_{3})=n_{3}(s)-1,&n_{1}(s_{3})>0,&n_{3}(s_{3})<0,\\ n_{2}(s_{4})=n_{3}(s),&n_{1}(s_{4})>0,&n_{3}(s_{4})<0,\\ n_{3}(s_{5})=n_{3}(s)-1,&n_{1}(s_{5})>0,&n_{2}(s_{5})>0.\end{matrix}\right$.
    Therefore, to address the URF data format nuisance, the position embedding can consider the relative distances to some tokens.
D.2 Experiment Details
D.2.1 Unaligned Copy
An instance of the unaligned copy task is like
$\mathtt{b\,x_{0}\,\dots\,x_{n-1}\,=\,x_{0}\,\dots\,x_{n-1}\,e}.$
The model is given the input $\mathtt{bx_{0}\dots x_{n-1}=}$ and expected to output the copy of $\mathtt{x_{0}\dots x_{n-1}}$. We use $\mathtt{b}$” and $\mathtt{e}$” instead of $\mathtt{[BOS]}$” and $\mathtt{[EOS]}$” for a simpler implementation with the GPT-2 tokenizer.
We sample 2000 $n$-length instances for each $n=1,\dots,5$ as the training data. In the evaluation, we examine the learned models on instances of length 1101-10. We train GPT-2 with key-only RPE and RPE-Square, respectively. The model is trained by AdamW with the cosine scheduler, where the initial learning rate is $0.0005$, the weight decay is $1.0$, the warmup ratio is $0.05$, the gradient accumulation step is $2$, and the per-device training batch size is $256$. We set the training steps to $10000$ but early stop at step $1000$ as the model with RPE-Square has achieved nearly perfect length generalization while the model with RPE shows almost no length generalization then.
From the perspective of LDHD generalization, the latent variable corresponding to the input $\mathtt{bx_{0}\dots x_{n-1}=x_{0}\dots x_{k}}$ is
$\left[(x_{0},x_{0}),\dots,(x_{k},x_{k}),(x_{k+1},*),\dots,(x_{N},x_{N})\right].$
[…]
D.2.2 URF Addition
For the URF $n$-addition training data, we first sample the lengths of two addends uniformly from $\{1,\dots,n\}\times\{1,\dots,n\}$. For two addends of lengths $(n_{1},n_{2})$, we then samples from $10^{n_{1}-1},10^{n_{1}}-1]\times[10^{n_{2}-1},10^{n_{2}}-1]$ (If $n_{i}=0$, then the corresponding sample interval is $[0,9]$). This is to guarantee the addends are length-uniform. For the addition $x_{n_{1}-1}\dots x_{0}+y_{n_{2}-1}\dots y_{0}=z_{n_{3}}z_{n_{3}-1}\dots z_{0}$, the training instance is
@@ D.2.2 URF Addition
Here, we add spaces between the characters to ensure each is tokenized separately. In our experiments, we train with $10000$ URF $4$-addition samples.
We choose GPT-2 with key-only position embeddings as our model. For the RPE and RPE-Square settings, we augment the GPT-2 model with RPE and RPE-Square, respectively. The implementation is adapted from HuggingFace (Wolf et al., 2020).
We train the models by AdamW, with the initial learning rate $0.0005$, the weight decay $1.0$, and the cosine scheduler. The warmup ratio is $0.05$. The gradient accumulation step is $2$. The per-device training batch size is $128$. We train the models for $20000$ steps and $200000$ steps. The experiments are run on a server with Ubuntu. The models are trained on two NVIDIA GeForce RTX 3090 GPUs.
Appendix E Additional Experiments
E.1 Additional Evaluations of RPE-Square
We present more evaluation results of RPE-Square by comparing it against more different position embeddings in additional tasks. Concretely, we consider three extra tasks: ParityCoT, Multiplication (1 * N), and Division (1 * N). Across all tasks (including Addition and Copy), we compare RPE-Square with five position embeddings: RPE (Shaw et al., 2018), RoPE (Su et al., 2024), NoPE (Kazemnejad et al., 2024), ALiBi (Press et al., 2021), and Abacus (McLeish et al., 2024). We keep the hyperparameter setting for the addition position embeddings in Addition and Copy: the model is trained by AdamW with the cosine scheduler, the initial learning rate is $0.0005$, the weight decay is $1.0$, the warmup ratio is $0.05$, the gradient accumulation step is $2$, and the per-device training batch size is $256$. In the three extra tasks, we choose the initial learning rate as $0.0005$ and keep the other hyperparameters the same.
Figures 3-7 presents the experimental results. Both RPE-Square and Abacus achieve non-trivial length generalization performance in the five tasks, and the overall evaluation results of RPE-Square are better than those of Abacus. The other four position embeddings do not achieve good length generalization performance. The reason, from our LDHD perspective, is that RPE-Square and Abacus succeed in handling the unaligned data format, while the others do not.
Figure 3: Addition. The models are trained on URF 4-addition and tested on URF 5-addition
(a) RPE-Square
(b) RPE
(c) RoPE
(d) NoPE
(e) ALiBi
(f) Abacus
Figure 4: Copy. The models are trained on scales 1-5 and tested on scales 1-10.
(a) RPE-Square
(b) RPE
(c) RoPE
(d) NoPE
(e) ALiBi
(f) Abacus
Figure 5: ParityCoT. The models are trained on scales 1-5 and tested on scales 1-10. An instance is like $\mathtt{[BOS]x_{0}\dots x_{n-1}=y_{0}\dots y_{n-1}[EOS]}$”, where $y_{0}=x_{0}$ and $y_{k}=x_{k}\oplus y_{k-1}$ for $k=1,\dots,n-1$.
(a) RPE-Square
(b) RPE
(c) RoPE
(d) NoPE
(e) ALiBi
(f) Abacus
Figure 6: Multiplication (1N1*N). The models are trained on scales 1-5 and tested on scales 1-10. An instance is like $\mathtt{[BOS]x*y_{0}\dots y_{n-1}=z_{0}\dots z_{n}[EOS]}$”.
(a) RPE-Square
(b) RPE
(c) RoPE
(d) NoPE
(e) ALiBi
(f) Abacus
Figure 7: Division ($N/1$). The models are trained on scales 1-5 and tested on scales 1-10. An instance is like $\mathtt{[BOS]x\setminus y_{n-1}\dots y_{0}=z_{n-1}\dots z_{0}[EOS]}$”, where $z=y//x$.
E.2 Additional Application of the Position Embedding Design Principle
In this subsection, we present another application of the position embedding design principle besides RPE-Square. We consider a task called AdditionMod10, which computes the sum of the addends modulo 10, in the unaligned format. An instance takes the form
$\mathtt{[BOS]x_{0}\dots x_{n-1}+y_{0}\dots y_{n-1}=z[EOS]},$
where $z=(x+y)\bmod 10$. Due to the modulo operation, the output depends solely on the first digits of the addends. This means the positional embedding need to capture the absolute relative distances to specific tokens (namely, “[BOS]” and “+”). The “absolute value” handles the LDHD generalization in the latent space, and “the relative distances to some special tokens” deals with the unaligned data format. Following the same design principle used in RPE-Square, we introduce a new position embedding called RPE-Absolute to encode “the absolute value of the relative distances to some special tokens”. The expression of $\text{RPE-Absolute}_{i,j}$ is
$\sum_{1\leq k\leq i}\frac{\exp\left((W_{Q}x_{i})^{\intercal}(W_{K}x_{k})\right)}{\sum_{1\leq k^{\prime}\leq i}\exp\left((W_{Q}x_{i})^{\intercal}(W_{K}x_{k^{\prime}})\right)}R_{i-k},$
where $W_{K},W_{Q},R$ are the learnable parameters.
The results are shown in Figure 8. We compare GPT-2 with RPE-Absolute against GPT-2 with RPE. The training hyperparameters are identical to those in the experiments of RPE-Square with the initial learning rate $0.0001$. The results demonstrate that RPE-Absolute achieves length generalization, while RPE only generalizes within the training lengths. This finding further supports the effectiveness of our position embedding design principle.
(a) RPE-Absolute
(b) RPE
Figure 8: AdditionMod10. The models are trained on scales 1-5 and tested on scales 1-10.